Suppose a 10-year, 10 percent, semiannual coupon bond with a par value of R1 000

is currently selling for R1 135.90, producing a nominal yield to maturity of 8 percent.

However, the bond can be called after 5 years for a price of R1 050.

i) (3 points) What is the bond’s nominal yield to call?

ii) (2 points) If you bought this bond, do you think you would be more likely to earn the

YTM or the YTC? Why?


1
Expert's answer
2020-09-06T17:40:37-0400

i)

Yield to call =(Coupon payment +(Call value-Current price)/Years to call)/(Call value+Current price)/2


Where


"\frac{Sem-annual interest}{coupon}""=\frac{10\%}{2}\times1000=50"


"n=5 years\times2=10" periods to call


Price=PV=-1135.90


The yield to call"=\frac{50+(1050-1135.90)/10}{(1050+1135.90)/2}"


Call price=R1050


Yield to call(semi-annual)"=\frac{4141}{1093}=0.0379"


"0.0379\times100=3.79\%"


Annual yield to call"=3.79\%\times2"


"=7.5\%"


ii)

I would be more likely to earn YTC because the coupons bond its rate is 10% and the YTC is 7.53% i can sell new bonds as well as raise money.